OpenStudy (sanchez9457):

(t^2)y''-4ty'+6y=0 t>0 y1(t)=t^2 How do i find the 2nd solution using the method of reduction of order?

OpenStudy (dumbcow):

here is a good reference: http://tutorial.math.lamar.edu/Classes/DE/ReductionofOrder.aspx

OpenStudy (dumbcow):

\[y_2(t) = vt^{2}\] \[y'_2(t) = v' t^{2} + 2vt\] \[y''_2(t) = v''t^{2}+4v' t+2v\] plug that into diff equ simplify (a lot of terms cancel) \[v''t^{4}=0\] \[v'' = 0\] \[v = c_1 t\] so \[y_2(t) = c_1t^{3}\] and the general solution is: \[y(t) = c_1t^{3} +c_2t^{2}\]

OpenStudy (loser66):

yours is \[t^2y"-4ty'+6y =0\] in which, P(t)= 4t. Let \(y_2=Vy_1\) So, the formula for reduction of order is \[y_1V"+(2y_1'+P(t)y_1)V'=0\] Plug them in, you have \[t^2 V"+(2t^2 +4t*t^2)V'=0\\t^2V"+(2t^2 +4t^3)V'=0\] Let W=V' (*) so, yours is \[t^2 W' +(2t^2+4t^3)W=0\\ t^2W' = -(2t^2+4t^3)W\\\frac{W'}{W}=-\frac{2t^2+4t^3}{t^2}=-4t-2\] Now \[\frac{dW}{W}=(-4t-2)dt\]integral both sides, lnW = -2t^2 -2t \[W = e^{(-2t^2 -2t)}\text{Plug back to W = V' so,} V'=e^{(-2t^2-2t)}\] Again, take integral both sides to get V, after you have V, plug back to\(y_2 = V y_1\) to get the second solution. That's what my prof taught me. I don't know why it doesn't look like what dumbcow did.